OpenStudy (anonymous):

find x: y= x^2 +6x +15

OpenStudy (anonymous):

\[y=x^2+6x+9+6=\left( x+3 \right)^2+6,\left( x+3 \right)^2=y-6\] \[x+3=\pm \sqrt{y-6},x=-3 \pm \sqrt{y-6}\]